the length of rectangle is 8 cm more than its width.if its perimeter is 56 cm ,find its length and area?
Answer by macston(5194)

W=width; L=length=W+8; P=2(W+L)=2(W+(W+8))
2(W+(W+8))=56 cm divide each side by 2
W+W+8=28 subtract 8 from each side
2W=20 divide each side by 2
W=10 cm
L=W+8=10+8=18 cm ANSWER 1 Length is 18 cm
Area=L x W= 18 cm x 10 cm=180 cm^2 ANSWER 2 Area is 180 cm^2
CHECK
P=2(L+W)
56 cm=2(18 cm+10cm)
56 cm=2(28 cm)
56 cm=56 cm